Circular No. 6149

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) Phone 617-495-7244/7440/7444 TWX 710-320-6842 ASTROGRAM CAM

MARSDEN@CFA.HARVARD.EDU or GREEN@CFA.HARVARD.EDU (science)

JUPITER I (IO) J. R. Spencer and J. E. B. Spencer, Lowell Observatory; and D. Griep,

NASA Infrared Telescope Facility (IRTF), write: “The very bright volcanic hot spot seen on Mar. 2 (cf. IAUC 6143) was seen again in 3.5- and 3.8μm disk-resolved images of Io between Mar. 9.65 and 9.79 UT, but its brightness had decreased severalfold, so that it was only a small fraction of the brightness of the sunlit disk. Further analysis of the Mar. 2 images has yielded a location of 95^{o} west, 45^{o} south for the hot spot, with a current uncertainty of ∼ ±15^{o}; it was as bright as the sunlit disk at 4.8 μm on Mar. 2, and low-precision 4.8-μm photometry of Io on Feb. 28.85 suggests that the spot was bright then also. All observations were made with the NSFCAM 1- to 5-μm camera at the IRTF on Mauna Kea.”

SUPERNOVA 1995E IN NGC 2441 Further CCD photometry (cf. IAUC 6144): Feb. 14.084 UT, 17: ± 0.5

(D. Lane, M. Blake, P. Gray, and B. Werstiuk, Saint Mary’s University, Halifax; unfiltered pre-discovery image; comparison stars GSC 4369.0725 and 4369.0881); 28.215, R = 15.5 ± 0.1 (Lane et al.); Mar. 7.91, V = 16.37 ± 0.03 (H. Mikuž, Ljubljana).

SUPERNOVA 1995D IN NGC 2962 Further CCD photometry (cf. IAUC 6144): Feb. 28.248 UT, R = 13.8

± 0.1 (Lane et al.); Mar. 7.94, V = 14.08 ± 0.02, B − V = +0.57 ± 0.04 (H. Mikuž, Ljubljana).

Corrigendum. On IAUC 6144, line 6, for Feb. 19.07 read Feb. 18.07

SUPERNOVA 1995H IN NGC 3526 The position reported by R. H. McNaught on IAUC 6146 was from a

CCD image obtained on Mar. 7.71 UT by McNaught and D. J. Asher with the 1.0-m reflector at Siding Spring; a magnitude of V = 15.7 was derived from that image.

COMET 117P/HELIN-ROMAN-ALU 1 CCD photometry by S. M. Larson and C. W. Hergenrother with the

Steward Observatory’s 2.3-m reflector on Jan. 29.48 UT gave R = 19.7. Three 120-s coadded frames showed a coma 8′′ in diameter and a 35′′ tail.

1995 March 13 Daniel W. E. Green